# Euler-Heisenberg Black Holes in Einsteinian Cubic Gravity

### 10 Feb 2024 v1

@@ I Introduction
Recently, a different cubic theory of gravity in four dimensions has been proposed, which contributes to the dynamics of the field equations without a matter coupling. This construction is base on Lovelock’s quadratic densities and the following desired properties i) General Relativity is recovered by considering the coupling constants $\beta^{(2)}=0=\beta^{(3)}$, ii) on maximally symmetric backgrounds, the linearized theory (2) propagates the same degrees of freedom as linearized Einstein gravity, and iii) the field equations derived from (2) admit Schwarzschild-like solutions, which are characterized by a single metric function [^2] By Schwarzschild-like solutions we mean static spherically symmetric metrics that satisfy the condition $g_{tt}g_{rr}=-1$ while the rest of the components correspond to the line element on a surface of constant scalar curvature and we focus on spherical topology.. To quadratic order, the most general polynomial $\mathcal{R}^{(2)}$ satisfying these conditions is the Gauss-Bonnet term, $\chi_{4}$, which is not dynamic in 4D; while to cubic order the polynomial $\mathcal{R}^{(3)}$ is written as followsKubiznak
@@
As mentioned above, effective field theories describe the dynamics of low energies only up to a certain energy scale $\mathcal{M}$. In the case of gravitational effective theories, when they exceed this energy scale, they exhibit Ostrogradski instabilities. For the case of ECG theory, these unstable modes have been found while considering perturbations over static and spherically symmetric solutions DeFelice , though these modes have an associated energy scale above the regime of validity of the EFT JimenezCano . In detail, the cubic terms ($\beta\Pi$) of the gravitational Lagrangian (4) are suppressed with respect to the Einstein-Hilbert term by $1/\mathcal{M}^{2}$ while the coupling constant $\beta$ is related to the cutoff scale as $\mathcal{M}=1/\beta^{1/4}$. The existence condition for these instabilities is $r^{3}<<\beta r_{h}\mathcal{M}^{2}$ DeFelice , which in the case of $\mathcal{M}=1/\beta^{1/4}$ reduces to $r^{3}<<\beta^{1/2}r_{h}$. Therefore, this condition indicates that these cubic correction terms in (4) have to be smaller than the linear term outside the event horizon of a black hole in order to avoid pathologies [^3] Note that our notation is equivalent to that in DeFelice when $M=\mathcal{M}$ and $\alpha=\beta$.. In the present work, we consider our solutions are valid in the regime where $|\beta\Pi|<<|R|$, i.e. in the regime where the EFT is valid, hence avoiding instabilities. This approach would exclude all non-perturbative effects that can generate corrections to the Einstein-Hilbert action, as well as quantum effects that might solve the presence of ghosts Hawking .
@@ II Action and Equations
Before describing our search for solutions to the field equations (8)-(10), it is interesting to review some of the results found in the literature about black holes found within these two EFTs, which have been mostly studied separately. For example, the search of black hole solutions in General Relativity (GR) coupled to the Euler-Heisenberg electrodynamics were found by Yajima and Tamaki for magnetic charges only Yajima , while the equivalent electric solutions were discovered in MagosBreton ; AmaroMacias , where the authors adopted the Plebanski $\mathscr{H(P)}$ formulation. Efforts to include rotation were also pursued, as in NoraRot . Moreover, different properties of some of these solutions have been analyzed, such as the geodesic structureAmaroMacias , thermodynamic stabilityMagosBreton and the shadow contourAllahyari . Even beyond the usual GR structure, EH electrodynamics has also been studied in the context of Modified GravityStefanov ; GuerreroRubiera . On the other hand, black holes with an ECG correction term which generalised the Schwarzchild and Reissner-Norsdtröm metrics were described inBuenoCano2 ; HennigarMann , meanwhile slowly and extremal rotating black holes were obtained in AdairBuenoCano ; CanoPeriniguez , respectively. Finally, there are other interesting physical applications of these cubic curvature corrections beyond the common black hole solutions, such as in the inflationary epoch the universeArciniega1 ; Arciniega2 ; Arciniega3 , holographic studies BuenoCanoRuiperez ; EdelsteinGrandi , observational deviations from GRPoshteh1 ; Poshteh2 , and other compact object such as NUT NUTs and wormholesMehdizadeh solutions.
@@ IV.4 Gravitational lensing in the weak field limit
One of the first effects used to test the theory of General Relativity was gravitational lensing, which consists of the deviation of the trajectory of light when passing near a massive object. When considering the weak field limit, the lensing effect around our black holes occurs when massless particles pass far away from the photosphere and are slightly deflected from their trajectory. In this case, the deflection angles are assumed to be small, thus can be determined by the Gaussian curvature and the Gauss-Bonnet theorem for the optical metric using the method developed by Gibbons and WernerGibbonsWL (for further details see example OkyayWL ).

### 21 Feb 2024 v2

@@ I Introduction
Recently, a different cubic theory of gravity in four dimensions has been proposed, which contributes to the dynamics of the field equations without a matter coupling. This construction is base on Lovelock’s quadratic densities and the following desired properties i) General Relativity is recovered by considering the coupling constants $\beta^{(2)}=0=\beta^{(3)}$, ii) on maximally symmetric backgrounds, the linearized theory (2) propagates the same degrees of freedom as linearized Einstein gravity, and iii) the field equations derived from (2) admit Schwarzschild-like solutions, which are characterized by a single metric function [^2] By Schwarzschild-like solutions we mean static spherically symmetric metrics that satisfy the condition $g_{tt}g_{rr}=-1$ while the rest of the components correspond to the line element on a surface of constant scalar curvature and we focus on spherical topology, see eq. (14).. To quadratic order, the most general polynomial $\mathcal{R}^{(2)}$ satisfying these conditions is the Gauss-Bonnet term, $\chi_{4}$, which is not dynamic in 4D; while to cubic order the polynomial $\mathcal{R}^{(3)}$ is written as followsKubiznak
@@
As mentioned above, effective field theories describe the dynamics of low energies only up to a certain energy scale $\mathcal{M}$. In the case of gravitational effective theories, when they exceed this energy scale, they exhibit Ostrogradski instabilities. For the case of ECG theory, these unstable modes have been found while considering perturbations over static and spherically symmetric solutions DeFelice , though these modes have an associated energy scale above the regime of validity of the EFT JimenezCano . In detail, the cubic terms ($\beta\Pi$) of the gravitational Lagrangian (4) are suppressed with respect to the Einstein-Hilbert term by $1/\mathcal{M}^{2}$ while the coupling constant $\beta$ is related to the cutoff scale as $\mathcal{M}=1/\beta^{1/4}$. The existence condition for these instabilities is $r^{3}<<\beta r_{h}\mathcal{M}^{2}$ DeFelice , which in the case of $\mathcal{M}=1/\beta^{1/4}$ reduces to $r^{3}<<\beta^{1/2}r_{h}$. Therefore, this condition indicates that these cubic correction terms in (4) have to be smaller than the linear term outside the event horizon of a black hole in order to avoid pathologies [^3] Note that our notation is equivalent to that in DeFelice when $M=\mathcal{M}$ and $\alpha=\beta$.. In the present work, we consider our solutions are valid in the regime where $|\beta\Pi|<<|R|$, i.e. in the regime where the EFT is valid, hence avoiding instabilities. This approach would exclude all non-perturbative effects that can generate corrections to the Einstein-Hilbert action, as well as quantum effects that might solve the presence of ghosts Hawking  [^4] For an alternative approach to study quantum corrections, applied to the Schwarzschild metric, under a gravitational EFT see Battista .
@@ II Action and Equations
Before describing our search for solutions to the field equations (8)-(10), it is interesting to review some of the results found in the literature about black holes found within these two EFTs, which have been mostly studied separately. For example, the search of black hole solutions in General Relativity (GR) coupled to the Euler-Heisenberg electrodynamics were found by Yajima and Tamaki for magnetic charges only Yajima , while the equivalent electric solutions were discovered in MagosBreton ; AmaroMacias , where the authors adopted the Plebanski $\mathscr{H(P)}$ formulation. Efforts to include rotation were also pursued, as in NoraRot . Moreover, different properties of some of these solutions have been analyzed, such as the geodesic structureAmaroMacias , thermodynamic stabilityMagosBreton and the shadow contourAllahyari . Even beyond the usual GR structure, EH electrodynamics has also been studied in the context of Modified GravityStefanov ; GuerreroRubiera . On the other hand, black holes with an ECG correction term which generalised the Schwarzchild and Reissner-Norsdtröm metrics were described inBuenoCano2 ; HennigarMann as well as approximate analytical solutions using fast-converging continued-fraction expansions Sajadi . Moreover, the authors in AdairBuenoCano and CanoPeriniguez , respectively, obtained slowly and extremal rotating black holes with these cubic gravitational models. Finally, there are other interesting physical applications of these cubic curvature corrections beyond the common black hole solutions, such as in the inflationary epoch the universeArciniega1 ; Arciniega2 ; Arciniega3 , holographic studies BuenoCanoRuiperez ; EdelsteinGrandi , observational deviations from GRPoshteh1 ; Poshteh2 , and other compact object such as NUT NUTs and wormholesMehdizadeh solutions.
@@ IV.4 Gravitational lensing in the weak field limit
One of the first effects used to test the theory of General Relativity was gravitational lensing, which consists of the deviation of the trajectory of light when passing near a massive object. Recently, these lensing effects has been used to constrain the compactness and distortion parameters of gravitational lenses; e.g. when applied to the supermassive black holes $*{}^{*}$ and $*{}^{*}$ in Virbhadra1 ; Virbhadra2 . When considering the weak field limit, the lensing effect around our black holes occurs when massless particles pass far away from the photosphere and are slightly deflected from their trajectory. In this case, the deflection angles are assumed to be small, thus can be determined by the Gaussian curvature and the Gauss-Bonnet theorem for the optical metric using the method developed by Gibbons and WernerGibbonsWL (for further details see example OkyayWL ).